# Analysis of Hessian Scaling for Local and Global Costs in Variational Quantum Algorithms

### v1

@@ Abstract
Barren plateaus are typically characterized by vanishing gradients, yet the feasibility of curvature-based optimization fundamentally relies on the statistical resolvability of the Hessian matrix. In this work, we quantify the entrywise resolvability of the Hessian for Variational Quantum Algorithms at random initialization. By leveraging exact second-order parameter-shift rules, we derive a structural representation that reduces the variance of Hessian entries to a finite covariance quadratic form of shifted cost evaluations. This framework reveals two distinct scaling regimes that govern the sample complexity required to resolve Hessian entries against shot noise. For global objectives, we prove that Hessian variances are exponentially suppressed, implying that the number of measurement shots must scale as $O(e^{\alpha n})$ with the number of qubits $n$ to maintain a constant signal-to-noise ratio. In contrast, for termwise local objectives in bounded-depth circuits, the variance decay is polynomial and explicitly controlled by the backward lightcone growth on the interaction graph, ensuring that curvature information remains statistically accessible with $O(\mathrm{poly}(n))$ shots. Extensive numerical experiments across varying system sizes and circuit depths demonstrate these theoretical bounds and the associated sampling costs. Our results provide a rigorous criterion for the computational tractability of second-order methods at initialization.
@@ 1 Introduction
Variational quantum algorithms (VQAs) have emerged as a dominant paradigm for the noisy intermediate-scale quantum (NISQ) era, formulating problems as stochastic optimization over parameterized quantum circuits [cerezo2021variational, kandala2017hardware, Peruzzo2014, jin2024quantum, hu2024quantum, an2023linear, zhu2025quantum]. However, their scalability is fundamentally constrained by the geometry of the optimization landscape. Beyond the general NP-hardness of training [PhysRevLett.127.120502], a central obstruction is the barren plateau phenomenon [mcclean2018barren, larocca2025barren]: for broad classes of ansatz families, gradient variances decay exponentially with system size, rendering first-order optimization ineffective at random initialization [Grant2019initialization, bharti2022noisy].
@@
The mechanisms driving this first-order information loss are now well understood. Theoretical analyses link vanishing gradients to factors such as excessive entanglement [marrero2021entanglement], the globality of the cost operator [cerezo2021cost], and hardware noise [wang2021noise]. Geometrically, these landscapes are characterized by narrow gorges [arrasmith2022equivalence] or a proliferation of traps [PhysRevA.111.012441]. In response, a diverse array of diagnostic [larocca2022diagnosing] and mitigation strategies has been developed, encompassing architectural designs [PhysRevX.11.041011], specialized initialization techniques [Grant2019initialization, meng2026trainability], layerwise training [Skolik2021], and controls on ansatz expressibility [holmes2022expressibility].
@@
Despite this progress, a systematic second-order theory for landscape curvature remains underdeveloped. Curvature-aware strategies, including Newton-type updates and the quantum natural gradient [stokes2020quantum], are staples of scientific computing. However, their efficacy in the quantum regime depends entirely on the statistical accessibility of curvature information as the system scales. While prior work indicates that higher-order derivatives can also vanish [cerezo2021higher] and empirical studies have explored specific applications [sen2022hessian], existing theoretical analyses largely focus on global operator norms or specific backpropagation schemes [Bowles2025BackpropScalingPQC]. These approaches overlook the entrywise statistical resolvability of the Hessian at random initialization, a property essential for practical optimization where estimators are constructed from finite measurement budgets. Consequently, a fundamental question arises: when first-order gradients vanish, does second-order information remain statistically accessible, and what precise scaling laws govern its variance?
@@
In this work, we bridge this gap by quantifying the initialization scaling of Hessian-entry fluctuations. Focusing on hardware-efficient architectures with bounded-depth locality, we build our analysis upon two technical pillars: the parameter-shift rule [bergholm2018pennylane, crooks2019gradients, wierichs2022general], which provides an exact framework for derivative evaluation, and the established role of cost function locality in mitigating first-order plateaus [cerezo2021cost, uvarov2021locality]. By leveraging exact second-order parameter-shift identities, we express each Hessian entry as a constant-size linear combination of shifted objective evaluations. This structural representation reduces the entrywise variance $\operatorname{Var}_{\rho}(H_{jk})$ to a finite covariance–quadratic form, allowing us to control the scaling explicitly via backward-lightcone geometry. Figure 1 schematically illustrates the resulting regimes.
@@
Figure 1: Higher-order barren plateaus: global vs. local objectives. Schematic landscapes and expected scaling of resolvable Hessian-entry variance with system size: global costs can exhibit exponential suppression at initialization, while local costs retain a larger (polynomially decaying) resolvable curvature scale.
Our analysis reveals two distinct scaling regimes at random initialization. For global observables, we prove that the Hessian inherits the objective’s exponential concentration, implying that resolving curvature against shot noise requires a prohibitive $O(e^{\alpha n})$ complexity. Conversely, for $k$-local objectives, we derive polynomial variance bounds explicitly governed by the ratio of the causal lightcone volume $V_{G}(k+2rL)$ to the system size $n$. This geometric scaling ensures that, provided the lightcone remains sub-extensive, curvature information remains statistically accessible with $O(\mathrm{poly}(n))$ measurements. We demonstrate this theoretical framework through numerical experiments that verify the predicted scaling laws, analyze the Hessian eigenspectra, and quantify the practical sampling complexity required for curvature estimation.
@@
The remainder of the paper is structured as follows. In § 2, we establish the theoretical framework, including architectural assumptions and the relevant parameter-shift representations. § 3 presents our main theoretical analysis, deriving the explicit variance bounds and scaling laws that distinguish global from local objectives. These analytical predictions are demonstrated in § 4 through extensive numerical experiments. Finally, we discuss practical implications and future directions in § 5. Detailed proofs and supplementary technical results are deferred to the appendices.
@@ 2 Theoretical Background
We establish the framework for VQAs needed to analyze higher-order derivatives. We define the circuit model, objective functions, and locality notions that control operator spreading, and we derive the parameter-shift identities for gradients and Hessian entries used throughout the paper. We also state the initialization assumptions and our definition of higher-order barren plateaus, which underpin the variance bounds in § 3.
@@ 2.1 Parameterized quantum circuits
We focus on hardware-efficient architectures of depth $L$ obtained by alternating fixed entangling layers and parameterized rotations. Specifically, each layer $\ell$ consists of a parameter-independent entangling unitary $V_{\ell}\in U(2^{n})$ followed by $m_{\ell}$ parametrized gates, so that
@@
with products ordered in increasing $\ell$ (as written) and $M=\sum_{\ell=1}^{L}m_{\ell}$. When convenient we index parameters by a single label $a\in\{1,\ldots,M\}$, with $a$ corresponding to a unique pair $(\ell,j)$ in (2).
@@
[Two-point-spectrum generators] In (2), each generator $G_{\ell,j}$ is Hermitian and supported on at most $r$ qubits, where $r$ is independent of $n$. Moreover, $G_{\ell,j}$ has a two-point spectrum $\mathrm{spec}(G_{\ell,j})=\{\pm\gamma\}$ for some $\gamma>0$. Without loss of generality we take $\gamma=\tfrac{1}{2}$, i.e.,
(3) $G_{\ell,j}^{2}=\frac{1}{4}I\qquad\text{equivalently}\qquad\mathrm{spec}(G_{\ell,j})=\{\pm\tfrac{1}{2}\}.$
This includes Pauli rotations $G_{\ell,j}=P_{\ell,j}/2$ with $P_{\ell,j}$ a Pauli string supported on at most $r$ qubits.
@@ 2.2 Cost functions and locality
An observable $O$ is $k$-local if it acts nontrivially on at most $k$ qubits, i.e., $|\operatorname{supp}(O)|\leq k$ for a constant $k$ independent of $n$. More generally, an objective $C(\bm{\theta})$ is $k$-local termwise if it can be written as $C(\bm{\theta})=\frac{1}{n}\sum_{i=1}^{n}\langle\psi(\bm{\theta})|\,O_{i}\,|\psi(\bm{\theta})\rangle$ with each $O_{i}$ being $k$-local. We call an observable $O$ (and the objective it induces via (4)) global if it has extensive support, i.e., $|\operatorname{supp}(O)|$ grows linearly with $n$.
@@
which acts nontrivially on all qubits. In contrast, local objectives are typically constructed from bounded-support terms, e.g., $Z_{i}$ or $Z_{i}Z_{i+1}$. For normalization one often considers averaged local costs, for instance
@@
which remains $1$-local termwise while producing an $O(1)$-scaled objective.
@@ 2.4 Trainability metrics at random initialization
We analyze gradient and Hessian fluctuations at random initialization. Unless otherwise stated, all expectations and variances are with respect to $\bm{\theta}\sim\rho$, where $\rho$ is the product measure of i.i.d. uniforms on $0,2\pi)$. We assume one parameter per gate instance in (2) (i.e., no parameter tying), so that $\bm{\theta}=\{\theta_{\ell,j}\}_{(\ell,j)\in\mathcal{I}}$ indexes all parametrized gates.
@@
[Observable normalization] The observable $O$ in (4) is Hermitian and satisfies $\|O\|\leq 1$. For termwise local objectives $C(\bm{\theta})=\frac{1}{n}\sum_{i=1}^{n}\langle\psi(\bm{\theta})|O_{i}|\\ \psi(\bm{\theta})\rangle$, we also assume $\|O_{i}\|\leq 1$.
@@ 2.4.1 Variance of Hessian entries
For any indices $(j,k)$, we quantify Hessian-entry fluctuations at initialization by
@@
We view $\operatorname{Var}_{\rho}(H_{jk})$ as a measure of the statistical resolvability of entrywise curvature under finite-budget derivative estimators.
@@
Definition 2.2 is entrywise and does not by itself imply spectral conditioning (e.g., bounds on extremal eigenvalues or condition numbers); small typical entries can still coexist with a few directions of non-negligible curvature. Nevertheless, entrywise second-moment control is the natural notion for finite-budget shift-based estimation and yields, for instance, $\mathbb{E}_{\rho}[\|H\|_{2}^{2}]\leq\mathbb{E}_{\rho}[\|H\|_{F}^{2}]=\sum_{j,k}\mathbb{E}_{\rho}[H_{jk}^{2}]$.
The following lemma reduces Hessian-entry variance to a covariance quadratic form over finitely many shifted objective evaluations, which is the starting point for our global/local scaling analysis. The proof is given in Appendix A.1.1.
@@
Under Assumption 2.1, the diagonal Hessian entry admits the second-order parameter-shift representation
(12) $H_{jj}=\frac{1}{4}\Big(C(\bm{\theta}+\pi\mathbf{e}_{j})-2C(\bm{\theta})+C(\bm{\theta}-\pi\mathbf{e}_{j})\Big).$
Define $C_{+}:=C(\bm{\theta}+\pi\mathbf{e}_{j})$, $C_{0}:=C(\bm{\theta})$, and $C_{-}:=C(\bm{\theta}-\pi\mathbf{e}_{j})$, and let $\Sigma\in\mathbb{R}^{3\times 3}$ be the covariance matrix with entries $\Sigma_{\alpha\beta}:=\operatorname{Cov}_{\rho}(C_{\alpha},C_{\beta})$ for $\alpha,\beta\in\{+,0,-\}$. Then
(13) $\operatorname{Var}_{\rho}(H_{jj})\;=\;a^{\top}\Sigma a,\qquad a:=\tfrac{1}{4}(1,-2,1)^{\top},$
and equivalently $\operatorname{Var}_{\rho}(H_{jj})$ can be written as an explicit linear combination of variances and covariances among $\{C_{+},C_{0},C_{-}\}$. An analogous covariance–quadratic representation holds for off-diagonal entries $H_{jk}$ using (2.3) (see Appendix A.1.2).
@@ 3 Asymptotic Variance Analysis
This section analyzes the asymptotic scaling of Hessian-entry variances. We first reduce $\operatorname{Var}_{\rho}(H_{jk})$ to a finite covariance–quadratic form over shifted cost evaluations, and then bound the resulting covariances using locality and lightcone structure for termwise local objectives, or a second-moment concentration condition for global objectives.
3.1 Variance representation via parameter shifts
Let $C(\bm{\theta})$ be the cost (4) and let $H_{jk}$ denote the Hessian entries given by (2.3)–(9). By Lemma 2.3, each $H_{jk}$ is a fixed linear combination of finitely many shifted costs, hence $\operatorname{Var}_{\rho}(H_{jk})$ reduces to a finite covariance–quadratic form.
For each $(j,k)$ there exist a finite shift set $\mathcal{S}_{jk}\subset\mathbb{R}^{M}$ and coefficients $\{a_{s}\}_{s\in\mathcal{S}_{jk}}$, depending only on whether $j=k$ or $j\neq k$, such that
(14) $\operatorname{Var}_{\rho}(H_{jk})=\operatorname{Var}_{\rho}\Big(\sum_{s\in\mathcal{S}_{jk}}a_{s}C(\bm{\theta}\oplus s)\Big)=\sum_{s,s^{\prime}\in\mathcal{S}_{jk}}a_{s}a_{s^{\prime}}\operatorname{Cov}_{\rho}\big(C(\bm{\theta}\oplus s),C(\bm{\theta}\oplus s^{\prime})\big),$
where $\oplus$ denotes componentwise addition modulo $2\pi$. For the two-shift rules (2.3)–(9) one has $|\mathcal{S}_{jk}|\leq 4$ for all $j,k$, and the coefficients $\{a_{s}\}$ are $O(1)$ constants independent of $n$. Accordingly, all constants in the bounds below may depend on the chosen shift rule but not on $n$. Eq. (14) will serve as the starting point for the variance bounds derived below.
3.2 Main result
This subsection derives explicit scaling bounds from (14) by bounding the covariances of the shifted costs. We begin with the global case and then turn to termwise $k$-local averaged objectives.

### v2

@@ Abstract
Barren plateaus in variational quantum algorithms are typically described by gradient concentration at random initialization. In contrast, rigorous results for the Hessian, even at the level of entry-wise variance, remain limited. In this work, we analyze the scaling of Hessian-entry variances at initialization. Using exact second-order parameter-shift identities, we write $H_{jk}$ as a constant-size linear combination of shifted cost evaluations, which reduces $\operatorname{Var}_{\rho}(H_{jk})$ to a finite-dimensional covariance–quadratic form. For global objectives, under an exponential concentration condition on the cost at initialization, $\operatorname{Var}_{\rho}(H_{jk})$ decays exponentially with the number of qubits $n$. For local averaged objectives in bounded-depth circuits, $\operatorname{Var}_{\rho}(H_{jk})$ admits polynomial bounds controlled by the growth of the backward lightcone on the interaction graph. As a consequence, the number of measurement shots required to estimate $H_{jk}$ to fixed accuracy inherits the same exponential (global) or polynomial (local) scaling. Extensive numerical experiments over system size, circuit depth, and interaction graphs validate the predicted variance scaling. Overall, the paper quantifies when Hessian entries can be resolved at initialization under finite sampling, providing a mathematically grounded basis for second-order information in variational optimization.
@@ 1 Introduction
Variational quantum algorithms (VQAs) have emerged as a dominant paradigm for the noisy intermediate-scale quantum (NISQ) era, formulating problems as stochastic optimization over parameterized quantum circuits [an2023linear, cerezo2021variational, hu2024quantum, jin2024quantum, kandala2017hardware, Peruzzo2014, zhu2025quantum]. However, their scalability is fundamentally constrained by the geometry of the optimization landscape. Beyond the general NP-hardness of training [PhysRevLett.127.120502], a central obstruction is the barren plateau phenomenon [larocca2025barren, mcclean2018barren]: for broad classes of ansatz families, gradient variances decay exponentially with system size, rendering first-order optimization ineffective at random initialization [bharti2022noisy, Grant2019initialization].
@@
The mechanisms driving this first-order information loss are now well understood. Theoretical analyses link vanishing gradients to factors such as excessive entanglement [marrero2021entanglement], the globality of the cost operator [cerezo2021cost], and hardware noise [wang2021noise]. Geometrically, these landscapes are characterized by narrow gorges [arrasmith2022equivalence] or a proliferation of traps [PhysRevA.111.012441]. In response, a diverse array of diagnostic [larocca2022diagnosing] and mitigation strategies has been developed, encompassing architectural designs [PhysRevX.11.041011], specialized initialization techniques [Grant2019initialization, meng2026trainability], layer-wise training [Skolik2021], and controls on ansatz expressibility [holmes2022expressibility].
@@
Despite this progress, rigorous quantitative results for Hessian structure and scaling at initialization remain limited. Second-order methods, such as Newton-type updates [meng2026trainability] and the quantum natural gradient [stokes2020quantum], are standard tools in scientific computing, but in the quantum setting their usefulness is constrained by how accurately Hessian information can be estimated as the system size grows. While prior work indicates that higher-order derivatives can also vanish [cerezo2021higher] and empirical studies have explored specific applications [sen2022hessian], existing theoretical analyses largely focus on global operator norms or specific backpropagation schemes [Bowles2025BackpropScalingPQC]. These approaches do not quantify the entry-wise statistical resolvability of the Hessian at random initialization, which is essential in practice where Hessian estimators are formed from finite-shot measurements. Consequently, a fundamental question arises: when first-order gradients vanish, does second-order information remain statistically accessible, and what precise scaling laws govern its variance?
@@
In this work, we bridge this gap by quantifying the initialization scaling of Hessian-entry variance at random initialization. Focusing on hardware-efficient architectures with bounded-depth locality, our analysis rests on two ingredients: (i) exact parameter-shift rules [bergholm2018pennylane, crooks2019gradients, wierichs2022general] for derivative evaluation, and (ii) the role of cost-function locality in mitigating first-order barren plateaus [cerezo2021cost, uvarov2021locality]. Using exact second-order parameter-shift identities, we express each Hessian entry as a constant-size linear combination of shifted objective evaluations. This representation rewrites the entrywise variance $\operatorname{Var}_{\rho}(H_{jk})$ as a finite covariance–quadratic form and makes its scaling explicitly controllable through backward-lightcone geometry.
@@
Figure 1: Higher-order barren plateaus: global vs. local objectives. Schematic landscapes and the expected scaling of Hessian-entry variance with system size: global costs exhibit exponential decay, whereas term-wise local costs exhibit polynomial decay.
Our analysis yields two distinct scaling regimes at random initialization. Fig. 1 schematically illustrates these regimes, which we refer to as higher-order barren plateaus. For global objectives, we show that Hessian-entry variances inherit the exponential concentration of the objective, implying an exponentially growing finite-shot cost to resolve individual Hessian entries as the system size $n$ increases. In contrast, for term-wise $k$-local objectives (sums of terms acting on at most $k$ qubits) implemented by bounded-depth circuits, we obtain polynomial variance bounds controlled by the growth of the backward lightcone on the interaction graph. When the lightcone remains sub-extensive, these bounds yield a polynomial resolution cost for Hessian-entry estimation at initialization. We support the theory with numerical experiments that verify the predicted scaling laws, probe the Hessian eigenspectra at initialization, and quantify the corresponding sampling costs implied by our resolvability criterion.
@@
The remainder of the paper is structured as follows. In § 2, we establish the theoretical framework, including architectural assumptions and the relevant parameter-shift representations. § 3 presents our main theoretical analysis, deriving the explicit variance bounds and scaling laws that distinguish global from local objectives. These analytical predictions are demonstrated in § 4 through extensive numerical experiments. Finally, we discuss practical implications and future directions in § 5. Detailed proofs and supplementary technical results are provided in the appendices.
@@ 2 Theoretical Background
We introduce the basic ingredients needed for the analysis of higher-order derivatives in VQAs. Specifically, the circuit model and objective functions, formalize locality notions that control operator spreading, and derive the parameter-shift identities for gradients and Hessian entries used throughout the paper. The initialization assumptions and the definition of higher-order barren plateaus are also stated here; these will be used to obtain the variance bounds in § 3.
@@ 2.1 Parameterized quantum circuits
We focus on hardware-efficient architectures of depth $L$ obtained by alternating fixed entangling layers and parameterized rotations. Specifically, each layer $\ell$ consists of a parameter-independent entangling unitary $V_{\ell}$ acting on $(\mathbb{C}^{2})^{\otimes n}$, followed by $m_{\ell}$ parametrized gates, so that
@@
with products ordered in increasing $\ell$ and $M=\sum_{\ell=1}^{L}m_{\ell}$. We assume that each parametrized gate in (2) carries its own independent parameter (i.e., no parameter sharing), so that the parameter vector can be written as ${\bm{\theta}}=\{\theta_{\ell,j}\}_{\ell=1,\dots,L;\,j=1,\dots,m_{\ell}}$.
@@
[Two-point spectrum generators] In (2), each generator $G_{\ell,j}$ is Hermitian and supported on at most $r$ qubits, where $r$ is independent of $n$. Moreover, $G_{\ell,j}$ has a two-point spectrum $\mathrm{spec}(G_{\ell,j})=\{\pm\gamma\}$ for some $\gamma>0$. Without loss of generality we take $\gamma=\tfrac{1}{2}$, i.e., $G_{\ell,j}^{2}=\frac{1}{4}I$. This includes Pauli rotations $G_{\ell,j}=P_{\ell,j}/2$ with $P_{\ell,j}$ a Pauli string supported on at most $r$ qubits.
@@ 2.2 Cost functions and locality
An observable $O$ is $k$-local if it acts non-trivially on at most $k$ qubits, i.e., $|\operatorname{supp}(O)|\leq k$ for a constant $k$ independent of $n$. More generally, an objective $C(\bm{\theta})$ is $k$-local term-wise if it can be written as $C(\bm{\theta})=\frac{1}{n}\sum_{i=1}^{n}\langle\psi(\bm{\theta})|\,O_{i}\,|\psi(\bm{\theta})\rangle$ with each $O_{i}$ being $k$-local. We call an observable $O$ (and the objective it induces via (3)) global if it has extensive support, i.e., $|\operatorname{supp}(O)|$ grows linearly with $n$.
@@
which acts non-trivially on all qubits. In contrast, local objectives are typically constructed from bounded-support terms, e.g., $Z_{i}$ or $Z_{i}Z_{i+1}$. For normalization one often considers averaged local costs, for instance
@@
which remains $1$-local term-wise while producing an $O(1)$-scaled objective.
@@ 2.4 Trainability metrics at random initialization
We study the initialization-time variance of gradient and Hessian entries. Unless otherwise stated, all expectations and variances are taken with respect to ${\bm{\theta}}\sim\rho$, where $\rho$ is the product measure of i.i.d. uniform random variables on $0,2\pi)$.
@@
[Observable normalization] The observable $O$ in (3) is Hermitian and satisfies $\|O\|\leq 1$. For term-wise local objectives $C(\bm{\theta})=\frac{1}{n}\sum_{i=1}^{n}\langle\psi(\bm{\theta})|O_{i}|\\ \psi(\bm{\theta})\rangle$, we also assume $\|O_{i}\|\leq 1$.
@@ 2.4.1 Variance of Hessian entries
For any indices $(j,k)$, we quantify the initialization-time variability of a Hessian entry by
@@
We interpret $\operatorname{Var}_{\rho}(H_{jk})$ as a proxy for the statistical resolvability of the entry $H_{jk}$ under finite-shot estimation: smaller variance indicates that substantially more samples are required to distinguish the entry from sampling noise at a prescribed accuracy.
@@
Definition 2.2 is entry-wise and does not by itself imply spectral conditioning (e.g., bounds on extremal eigenvalues or condition numbers); small typical entries can still coexist with a few directions of non-negligible curvature. Nevertheless, entry-wise second-moment control is the natural notion for finite-sample shift-based estimation and yields, for instance, $\mathbb{E}_{\rho}[\|H\|_{2}^{2}]\leq\mathbb{E}_{\rho}[\|H\|_{F}^{2}]=\sum_{j,k}\mathbb{E}_{\rho}[H_{jk}^{2}]$ (cf. Lemma 3.9).
The following lemma reduces Hessian-entry variance to a covariance–quadratic form over finitely many shifted objective evaluations, which is the starting point for our global/ local scaling analysis. The proof is given in Appendix A.1.1.
@@
Under Assumption 2.1, the diagonal Hessian entry admits the second-order parameter-shift representation (8). Define $C_{+}:=C(\bm{\theta}+\pi\mathbf{e}_{j})$, $C_{0}:=C(\bm{\theta})$, and $C_{-}:=C(\bm{\theta}-\pi\mathbf{e}_{j})$, and let $\Sigma\in\mathbb{R}^{3\times 3}$ be the covariance matrix with entries $\Sigma_{\alpha\beta}:=\operatorname{Cov}_{\rho}(C_{\alpha},C_{\beta})$ for $\alpha,\beta\in\{+,0,-\}$. Then
(11) $\operatorname{Var}_{\rho}(H_{jj})=w^{\top}\Sigma w,\qquad w:=\tfrac{1}{4}(1,-2,1)^{\top},$
and equivalently $\operatorname{Var}_{\rho}(H_{jj})$ can be written as an explicit linear combination of variances and covariances among $\{C_{+},C_{0},C_{-}\}$. An analogous covariance–quadratic representation holds for off-diagonal entries $H_{jk}$ using (2.3). More generally, whenever a derivative quantity admits a finite-shift representation $X(\bm{\theta})=\sum_{\ell=1}^{m}w_{\ell}\,C(\bm{\theta}\oplus s_{\ell})$ with $m<\infty$, one has $\operatorname{Var}_{\rho}(X)=\sum_{\ell,\ell^{\prime}=1}^{m}w_{\ell}w_{\ell^{\prime}}\operatorname{Cov}_{\rho}\!\big(C(\bm{\theta}\oplus s_{\ell}),C(\bm{\theta}\oplus s_{\ell^{\prime}})\big)$.
@@ 3 Asymptotic Variance Analysis
This section derives asymptotic bounds for the entry-wise Hessian variance at random initialization. We first express $\operatorname{Var}_{\rho}(H_{jk})$ as a finite covariance–quadratic form over a constant number of shifted cost evaluations. We then bound these covariances using locality and backward-lightcone structure for local objectives, and using a cost-concentration condition for global objectives.
3.1 Variance representation
Let $C(\bm{\theta})$ be the cost (3). The second-order parameter-shift rules (2.3)–(8) express each Hessian entry as a finite linear combination of shifted costs, and hence yield the covariance–quadratic form
(12) $\operatorname{Var}_{\rho}(H_{jk})=\sum_{s,s^{\prime}\in\mathcal{S}_{jk}}w_{s}w_{s^{\prime}}\operatorname{Cov}_{\rho}\big(C(\bm{\theta}\oplus s),C(\bm{\theta}\oplus s^{\prime})\big),$
where $\oplus$ denotes component-wise addition modulo $2\pi$. For the two-shift rules (2.3)–(8), one has $|\mathcal{S}_{jk}|\leq 4$ and $w_{s}=O(1)$ independent of $n$. Derived from Lemma 2.3, the explicit diagonal and off-diagonal shift rules are listed in Appendix A.1; (12) is the starting point for the variance bounds in § 3.2.
While (12) characterizes the initialization-induced variability of $H_{jk}$, practical estimation relies on finite-shot evaluations of the shifted costs. It is therefore natural to compare the shot-noise contribution to the estimator variance with the intrinsic scale $\operatorname{Var}_{\rho}(H_{jk})$ and to quantify the number of shots needed to separate the two. This motivates the following resolution criterion.
Definition 3.1 (Resolution criterion and shot requirement).
Let $\widehat{H}_{jk}$ be a finite-shot estimator of $H_{jk}$ constructed from shifted cost evaluations. For $\eta\in(0,1]$, we say that $H_{jk}$ is resolved at initialization if
$\mathbb{E}_{\bm{\theta}\sim\rho}\left[\operatorname{Var}_{\mathrm{sh}}(\widehat{H}_{jk}\mid\bm{\theta})\right]\leq\eta\operatorname{Var}_{\rho}(H_{jk}).$
The corresponding resolution shot requirement $N_{\mathrm{res}}$ is the smallest number of shots per shifted cost evaluation for which the above condition holds.
Algorithm 1 states the finite-shot parameter-shift estimator used throughout and in the experiments of § 4.
Algorithm 1 Finite-shot estimation of a Hessian entry via parameter-shift
1:Ansatz $U(\bm{\theta})$, objective $C(\bm{\theta})$, indices $(j,k)$, measurement shots NN
2:Unbiased estimate $\widehat{H}_{jk}$ of the Hessian entry $H_{jk}$
[…]
4:for all $s\in\mathcal{S}_{jk}$ do $\triangleright$ Loop over shift set $\mathcal{S}_{jk}$ with coeff. $w_{s}$ (cf. § 2.3)
5:  Prepare $|\psi(\bm{\theta}\oplus s)\rangle$ and estimate $C(\bm{\theta}\oplus s)=\langle O\rangle$ using NN shots
7:  $\widehat{H}_{jk}\leftarrow\widehat{H}_{jk}+w_{s}\widehat{C}(\bm{\theta}\oplus s)$
8:end for
9:return $\widehat{H}_{jk}$
3.2 Main results
We now derive explicit scaling bounds from (12) by bounding covariances between shifted costs. We treat global objectives first, where second-moment concentration transfers through the constant-size finite-shift rules, and then turn to term-wise $k$-local averaged objectives.